# ID-Pose: Sparse-view Camera Pose Estimation by Inverting Diffusion Models

### v1

@@ 4.2 Evaluation Results
Method $15^{\circ}$ $30^{\circ}$ PA $15^{\circ}$ $30^{\circ}$ PA $15^{\circ}$ $30^{\circ}$ PA $15^{\circ}$ $30^{\circ}$ PA
RelPose++ w/o Mask 5.56 14.20 8.64 4.12 12.35 7.20 3.70 12.72 8.27 4.41 12.61 9.08
RelPose++ 6.17 11.11 12.96 4.32 11.73 12.76 3.70 11.85 11.60 4.50 11.90 11.46
ID-Pose w/o Tri 48.77 56.79 45.06 48.77 59.88 44.65 48.77 60.25 44.57 49.65 61.38 45.06
ID-Pose 48.77 56.79 45.06 49.79 61.11 46.50 50.12 63.33 46.42 52.38 65.26 48.06
@@
Table 2: Results on the OmniObject3D testing samples. “$15^{\circ}$”: rotation accuracy by $15^{\circ}$. “$30^{\circ}$”: rotation accuracy by $30^{\circ}$. “PA”: camera position accuracy. The numbers are shown in 100%.
$n=2$ $n=4$ $n=6$ $n=8$
Method $15^{\circ}$ $30^{\circ}$ PA $15^{\circ}$ $30^{\circ}$ PA $15^{\circ}$ $30^{\circ}$ PA $15^{\circ}$ $30^{\circ}$ PA
RelPose++ w/o Mask 2.00 6.00 2.00 2.00 7.33 1.67 3.00 7.40 3.20 2.86 7.86 3.43
RelPose++ 2.00 6.00 1.00 2.00 7.00 1.67 4.00 7.60 4.00 3.43 7.29 3.86
ID-Pose w/o Tri 76.00 80.00 73.00 75.00 78.33 74.00 77.60 80.00 75.60 77.71 80.57 75.86
ID-Pose 76.00 80.00 73.00 77.00 79.33 74.00 78.00 81.40 75.60 78.71 82.14 77.14
Table 3: Results on the ABO testing samples.
We report evaluation results aggregated by the number of input views. The results on OmniObject3D and ABO samples are shown in Table 2 and Table 3, respectively. ID-Pose methods significantly outperform RelPose++ for all different $n$, which demonstrate advanced generalization capability. RelPose++ methods report much lower metrics compared to the testing results on the CO3D dataset [7]. The reason might be that the models of RelPose++ overfit to CO3D data and fail to generalize to out-of-distribution images. With two input images ($n=2$), ID-Pose w/o Tri and ID-Pose report the same results as they behave identically. With more input images ($n>2$), ID-Pose outperforms ID-Pose w/o Tri as extra image pairs are used to estimate poses. By adding input images, the performance of ID-Pose increase as more image pairs are used.
We visualize examples of estimated camera poses in the world coordinate system. We compare ID-Pose and Relpose++ with ground truth. Assume that the true camera pose of the reference image is known. For RelPose++, we transform all predicted poses with a matrix that aligns the first prediction to the reference. In Figure 2, we input four views of the “house_001” from OmniObject3D, where the front view acts as the reference. Compared with the ground truth, ID-Pose finds all three relative poses with small errors. RelPose++ accurately predicts the relative poses of the two side views, but makes a large prediction error on the opposite view. In Figure 3, we input four views of the “ornaments_001” from OmniObject3D, where the “left-face” view acts as the reference. This is a hard case for both ID-Pose and RelPose++. ID-Pose finds the approximate poses but with large errors compared to the ground truth. RelPose++ predicts poses with even larger errors. In Figure 4, we input four views of an ABO chair object. ID-Pose accurately finds the camera poses, while RelPose++ fails to distinguish the differences between these images. This failure happens to almost all the ABO testing samples. RelPose++ is probably hard to handle out-of-distribution images. This exposes the weakness of the fitting based methods, that the generalization capability is limited by the amount of annotated training data. ID-Pose leverages the model pre-trained on large-scale images, which demonstrates better generalization ability.
Figure 2: Visualization of camera poses with a sample of views from OmniObject3D “house_001”.
Figure 3: Visualization of camera poses with a sample of views from OmniObject3D “ornaments_001”.
Figure 4: Visualization of camera poses with a sample of views from an ABO chair.
4.3 Qualitative Results on In-the-wild Images
@@
We qualitatively show the results of ID-Pose on in-the-wild images. We find three objects in our daily workplace: a toy duck, a chair, and a foosball table. We use a smartphone to capture 4-6 images for each of them. We use the DIS method [13] to segment out the objects and clean the background of the images to white. We select $n$ ($n=2,3,4$) images as input views to estimate the relative poses. To visualize the results, we use the LumaAI APP [^1] https://lumalabs.ai/ to reconstruct 3D meshes of the objects. We then choose one image as the reference and manually find its camera pose using the 3D mesh. With the ground truth camera pose of the reference image, we transform the estimated poses of the target views into the world coordinate system. We show an example of the results of the chair in Figure 5. More results can be interactively visualized on the project website [^2] https://xt4d.github.io/id-pose/. The estimated poses cannot align the views to the 3D object perfectly. One reason is that the true cameras may not look at the same point in the world, and their “up” vectors can be varied. For in-the-wild images, pose estimation with the assumption of spherical coordinates will introduce additional errors. This is a weakness of ID-Pose.
@@
Figure 5: Visualization of estimated camera poses of sparsely captured views around a chair. The mesh is reconstructed using LumaAI APP. The view in red acts as the reference view, where the camera pose is manually found using the mesh.
5 Conculsion & Limitations
@@
In this paper, we propose ID-Pose which inverses the denoising diffusion process to estimate camera poses of sparse input views. ID-Pose only utilizes the pre-trained Zero-1-to-3 model [8] without additional training. Given two images, ID-Pose adds a noise on one image, and predicts the noise conditioned on the other image and a decision variable for the relative pose. The prediction error is used as the objective to find the optimal pose with the gradient descent method. ID-Pose can handle more than two images. It estimates each of the poses with additional pairs of images from triangular relationships. We conduct experiments using high-quality real-scanned 3D objects, where ID-Pose significantly outperforms state-of-the-art methods.
Limitations. The running time of ID-Pose is significantly higher than neural fitting based methods. It requires multiple steps of pose updating by inference through large diffusion networks. The average time for running 100 steps is about 30 seconds on an NVIDIA V100 GPU. This is time-consuming and the method needs efficiency improvement. Due to the limitation of Zero-1-to-3, ID-Pose assumes that all cameras look at the same point and have no rotation around optical direction. This will introduce estimation errors for in-the-wild images. To extend the capability, one possible solution is to fine-tune the novel view diffusion model with the input of 6DOF poses rather than spherical coordinates.

### v2

@@ 4.2 Evaluation Results
$15^{\circ}$ $30^{\circ}$ PA $15^{\circ}$ $30^{\circ}$ PA $15^{\circ}$ $30^{\circ}$ PA $15^{\circ}$ $30^{\circ}$ PA
NAVI Pairs 210 Pairs 630 Pairs 1050 Pairs 1470
COLMAP 25.24 26.67 0.00 16.35 16.51 0.16 12.67 12.76 0.10 11.09 11.16 0.41
$\vdash$ Converge # 81 + 0 of 210 32 + 62 of 210 11 + 85 of 210 11 + 87 of 210
RelPose++ FULL 38.10 63.33 21.43 36.03 62.38 21.90 35.71 61.62 22.95 36.67 62.93 22.52
RelPose++ MASK 25.24 51.43 17.14 25.08 49.52 15.87 24.10 49.71 15.90 22.45 48.37 13.88
ID-Pose Naive 48.57 81.43 42.38 53.33 79.52 44.29 52.10 79.24 44.10 53.20 79.18 43.88
ID-Pose TR 48.57 81.43 42.38 56.35 83.49 47.78 58.29 83.24 48.19 60.61 83.20 50.27
ID-Pose 48.57 81.43 42.38 57.30 85.56 48.25 60.19 86.76 50.48 62.52 87.62 51.84
OmniObject3D Pairs 162 Pairs 486 Pairs 810 Pairs 1134
COLMAP 12.35 15.43 0.00 8.23 9.88 0.00 8.27 9.63 0.37 8.73 9.88 0.88
$\vdash$ Converge # 53 + 0 of 162 20 + 43 of 162 16 + 51 of 162 19 + 51 of 162
RelPose++ FULL 12.35 30.25 8.64 9.47 27.16 7.20 9.88 27.16 8.27 10.93 30.51 9.08
RelPose++ MASK 16.05 38.27 12.96 13.37 34.36 12.76 12.72 33.83 11.60 12.17 33.33 11.46
ID-Pose Naive 70.99 86.42 62.35 64.81 80.66 54.73 64.44 80.25 55.06 63.49 79.81 55.11
ID-Pose TR 70.99 86.42 62.35 69.34 82.10 61.52 70.49 82.84 62.59 69.05 83.16 63.76
ID-Pose 70.99 86.42 62.35 67.70 81.07 59.47 71.23 84.07 64.20 70.11 84.13 64.02
ABO Pairs 100 Pairs 300 Pairs 500 Pairs 700
COLMAP 5.00 7.00 0.00 2.00 2.67 0.00 1.20 1.60 0.00 1.71 2.00 0.00
$\vdash$ Converge # 12 + 0 of 100 3 + 10 of 100 0 + 13 of 100 2 + 12 of 100
RelPose++ FULL 4.00 9.00 2.00 3.33 8.00 1.67 3.80 9.80 3.20 3.57 10.57 3.43
RelPose++ MASK 3.00 8.00 1.00 2.33 9.67 1.67 4.20 9.60 4.00 4.29 10.00 3.86
ID-Pose Naive 52.00 64.00 43.00 47.67 64.33 41.33 46.60 60.80 40.80 50.71 64.86 43.86
ID-Pose TR 52.00 64.00 43.00 53.00 64.67 45.33 52.80 62.00 46.80 57.86 65.71 50.29
ID-Pose 52.00 64.00 43.00 52.00 61.00 44.33 55.00 63.40 49.00 59.29 68.00 53.00
CO3D Pairs 180 Pairs 540 Pairs 900 Pairs 1260
COLMAP 20.00 25.56 0.00 19.44 21.67 0.93 22.44 24.11 2.00 23.89 26.51 3.25
$\vdash$ Converge # 86 + 0 of 180 48 + 54 of 180 54 + 57 of 180 60 + 60 of 180
RelPose++ FULL 83.33 92.22 47.22 84.63 94.44 49.44 85.78 94.11 51.11 86.83 94.52 50.24
RelPose++ MASK 81.67 90.00 46.67 82.41 91.30 49.63 81.44 91.22 49.33 82.22 91.59 47.38
ID-Pose Naive 39.44 64.44 32.22 42.22 66.30 32.59 42.00 69.67 32.78 41.51 67.30 31.83
ID-Pose TR 39.44 64.44 32.22 46.30 69.07 36.48 46.44 72.78 40.56 46.59 71.11 38.65
ID-Pose 39.44 64.44 32.22 47.22 70.56 37.96 46.56 73.56 39.89 48.49 72.70 39.29
@@
Table 2: Results on the testing samples of all datasets. “$15^{\circ}$”: rotation accuracy with $15^{\circ}$ threshold. “$30^{\circ}$”: rotation accuracy with $30^{\circ}$ threshold. “PA”: camera position accuracy. The values of the results are shown in 100%.
[…]
We report evaluation results aggregated by the number of input views $n$. The results on four datasets are shown in Table 2. We compare ID-Pose with the state-of-the-art RelPose++. According to the results on NAVI, OmniObject3D, and ABO samples, ID-Pose significantly outperforms RelPose++ for all different $n$. ID-Pose demonstrates advanced generalization capability on both casually captured photos and rendered images from arbitrary viewpoints. For CO3D samples, ID-Pose shows lower accuracy than RelPose++, which exhibits remarkable performances on this dataset. The reason might be that RelPose++ is trained with CO3D data and the testing samples share a similar distribution. However, as RelPose++ is based on fitting neural networks, the method could fail with out-of-distribution images and camera poses. In comparison, ID-Pose does not require training and generalizes well to a wide range of images. Note that RelPose++ FULL performs better than MASK on CO3D and NAVI, while performs worse on OmniObject3D and ABO. The reason is that the images of CO3D and NAVI include environmental background. This provides extra information which is learned by FULL to improve accuracy. The images of OmniObject3D and ABO has no background, which are out-of-distribution data for FULL. We find that COLMAP frequently fails to converge with sparse views, due to lack of reliable features from views with large appearance differences. We additionally report converge counts of running COLMAP. The counts includes 3 numbers A + B of C. A is the count of successful convergences with all $n$ images. B is the count of sub-success cases that the running with all images failed but there exists a success with fewer images. C is the total number of the samples. The successful count is higher for $n=2$, as joining two views is simple. With $n>2$, solving joint matches of multiple sparse views becomes difficult. We compare ID-Pose with the two ablated variants: Naive and TR. With $n=2$, they show the same results as there is no triangular relation. With $n>2$ , ID-Pose mostly outperforms the others, and TR outperforms the Naive. This demonstrates the effectiveness of using triangulation. There are some exceptions with $n=4$ for OmniObject3D and ABO, where TR performs the best. As these images are rendered from random viewpoints, there could be abnormal views that affect the exploration process to find promising initial poses.
We visualize examples of estimated camera poses in the world coordinate system, under the assumption that camera poses of anchor views are known. We compare ID-Pose with RelPose++ (best version). In Figure 2, we show two examples from NAVI and two examples from OmniObject3D, where the cameras are presented by 3D cones. The rotations and shifts between the estimated cones and the ground truth (grey color) can be observed. ID-Pose (blue color) accurately finds the camera poses, and surpasses the performance of RelPose++ (green color). As these objects are category-agnostic, ID-Pose shows better generalization ability due to leveraging the generative model pre-trained on large-scale images. Note that the bottom-right sample is a “dumbbell” which exhibits symmetry along its central axis. It is intractable to differentiate two symmetrically opposite views, e.g., image 1 and image 2, based on pure object appearance. However, ID-Pose conditions on lighting to find accurate camera poses. In Figure 3, we show two examples from ABO and two examples from CO3D. Given ABO samples, which are synthetic objects, ID-Pose yields accurate results as expected. But RelPose++ fails to predict the poses as it cannot generalize to synthetic content. Given CO3D samples, ID-Pose under-performs RelPose++ which might overfit to this particular data distribution. In general, ID-Pose based on iterative optimization finds more accurate poses than feed-forward prediction. We show some failed cases of ID-Pose. In the “chair” example with $n=4$, the pose for image 3 is incorrect. In the “bicycle” example with $n=6$, the poses of the image 2 and 5 are incorrect. We find that these prediction are axis-symmetric to the ground truth. This indicates that similar symmetric views can impact the performance of ID-Pose.
[…]
4.3 Noise Step Study
[…]
4.4 Using Self-captured Images
@@
We qualitatively show the results of ID-Pose on self-captured images. We find four objects in our daily workplace: a toy duck, a toy car, a chair, and a foosball table. We use a smartphone to capture 4-6 images for each of them, and use the DIS method [19] to segment out the objects with clean background. We select $n$ ($n=2,3,4$) images as input views to estimate the relative poses. To visualize the results, we use the LumaAI APP [^1] https://lumalabs.ai/ to reconstruct 3D meshes of the objects. We then choose one image as the anchor and manually find its camera pose using the 3D mesh. With the ground truth camera pose of the anchor image, we transform the estimated poses of the target views into the world coordinate system. We show an example of the results of the chair in Figure 5. More results are provided in our Project Page. We observe that the estimated poses cannot align the views to the 3D object perfectly. A reason is that the cameras may not look at the same point in the world, and their “up” vectors can be varied. Therefore, pose estimation with the assumption of spherical coordinates will introduce additional errors.
4.5 GPU Memory Usage & Running Time
We conduct experiments using a single NVIDIA V100 GPU 32G. Running the method requires loading a Zero-1-to-3 checkpoint, where the GPU memory usage is about 28G. The running time includes exploration time and refining time. For two images, the exploration time is about 1 minute. The refining takes about 30 seconds per 100 iterations. Running 600 iterations takes about 3 minutes.
Figure 4: The metrics with respect to noise step $t$.
@@ 4.5 GPU Memory Usage & Running Time
Figure 5: Visualization of the estimated camera poses of the captured images around a chair. The mesh is reconstructed using LumaAI APP. The view in red acts as the anchor view, where the camera pose is manually found using the mesh.
5 Conclusion & Limitations
@@
In this paper, we propose ID-Pose which inverses the denoising diffusion process to estimate camera poses of sparse input views. ID-Pose is a zero-shot method based on the pre-trained Zero-1-to-3 model [13]. To find the relative pose of two given images, ID-Pose uses the noise predictor, which conditions on one image and a hypothesis of the relative pose, to predict the noise added on another image. The prediction error is used as the objective function to find the optimal pose with the gradient descent method. ID-Pose can handle more than two images. It estimates each pose with additional pairs of images from triangular relations, where a joint optimization is applied to improve accuracy. We conduct extensive experiments on casually captured photos and rendered images with random viewpoints, where ID-Pose significantly outperforms state-of-the-art methods.
Limitations. The execution time of ID-Pose is considerably greater than that of feed-forward networks. The method requires multiple iterations of pose updating through extensive diffusion networks, which is a computationally costly process. ID-Pose may adopt a more efficient optimization approach to improve the time efficiency. Due to the limitation of Zero-1-to-3, ID-Pose operates under the assumption that all cameras look at the identical point and have no rotation around the optical axis. While the model can accommodate minor deviations, substantial shifts may result in estimation inaccuracies. To mitigate this issue, a potential solution is to fine-tune the novel view diffusion model with the input of 6DOF poses rather than spherical coordinates.